The equation of a parabola is y = ax2 + bx + c, where a, b and c are constants. The
x- and y-coordinates of a projectile launched from the origin as a function of time are
given by x = v0xt and y = v0yt − 1gt2, where v0x and v0y are the components of the 2
initial velocity.
(a) Eliminate t from these two equations and show that the path of a projectile is a
parabola and has the form y = ax2 + bx.
(b) What are the values of a, b and c for the projectile?
1
Expert's answer
2020-02-05T12:16:18-0500

From the conditions of the problem we have

"x=v_{0x}\cdot t" (1)

"y=v_{0y}\cdot t-g\cdot t^2" (2)

From (1) we write

"t=\frac{x}{v_{0x}}"

substitute in (2)

"y=v_{0y}\cdot \frac{x}{v_{0x}}-g\cdot (\frac{x}{v_{0x}})^2=x\cdot \frac{v_{0y}}{v_{0x}}-x^2\cdot \frac{g}{(v_{0x})^2}=-\frac{g}{(v_{0x})^2}\cdot x^2+\frac{v_{0y}}{v_{0x}}\cdot x" (3)

comparing equation (3) with equation "y=ax^2+bx+c"

get the expression

"a=-\frac{g}{(v_{0x})^2}"

"b=\frac{v_{0y}}{v_{0x}}"

"c=0"